OpenStudy (anonymous):

Write the equation of an ellipse with a major axis of length 8 and co-vertices (0,3) and (0,-3).

OpenStudy (anonymous):

Write an equation for an ellipse with center (1, –3), vertices (1, 2) and (1, –8), and co–vertices (4, –3) and (–2, –3). The distance between vertices (1,2),(1,-8) is sqrt[ (1-1)^2+(2+8)^2] = 10 2a=10 a=5 a^2=25 The distance between co-vertices (4,-3),(-2,-3) is sqrt[ (4+2)^2+(-3+3)^2] = 6 \[2b=6\] \[b=3\] \[b^2=9\] \[(x-1)^2/25 +(y+3)^2/9 = 1\]

OpenStudy (anonymous):

thats an example!

OpenStudy (anonymous):

 the equation, since it is centered at the origin is of this form: x^2/a^2 + y^2/b^2 = 1 can you take it from here?